OpenStudy (anonymous):

derivative of 1/sqrt(x^3 + 1) please show all work thanks in advance

mathslover (mathslover):

\[\large{\frac{d}{dx}{(\frac{1}{\sqrt{x^3+1}})}}\]

mathslover (mathslover):

Use the chain rule here

mathslover (mathslover):

do u know about chain rule @JayCT ?

OpenStudy (anonymous):

no can you help me pls

mathslover (mathslover):

Use the chain rule like this \[\large{\frac{d}{dx}(\frac{1}{\sqrt{x^3+1}})=\frac{d}{du}\frac{1}{\sqrt{u}}\frac{du}{dx}, \space \text{where} u = x^3 + 1\space \text{and} \space \frac{d}{du}{\frac{1}{\sqrt{u}} = \frac{-1}{2u^{\frac{3}{2}}}}}\]

OpenStudy (anonymous):

choose quotient rule or pure chain rule... chain rule requires a re-write to push the denominator to the numerator.

mathslover (mathslover):

so what we get now is : \[\large{\frac{\frac{d}{dx}(x^3+1)}{2(x^3+1)^{\frac{3}{2}}}}\] Differentiate the sum term by term : \[\large{\implies -\frac{\frac{d}{dx}x^3+\frac{d}{dx}(1)}{2(x^3+1)^{\frac{3}{2}}}}\] Since the deriviateive of a constant = 0

mathslover (mathslover):

therefore we get : \[\large{\frac{-\frac{d}{dx}x^3}{2(x^3+1)^{\frac{3}{2}}}}\] \[\large{\frac{-3x^2}{2(x^3+1)^{\frac{3}{2}}}}\]

mathslover (mathslover):

got it?

mathslover (mathslover):

This is much complicated... . You must be able to know what is : Chain rule and how to use that...

OpenStudy (anonymous):

as of now, no but ill try to go over and see can you help me with this last one (x^4 + x)^3 (x^2 - 1)^2 find the derivative pls

OpenStudy (anonymous):

overall... product rule... with chain rule inside

mathslover (mathslover):

ofcourse but ask that as a "new" question

OpenStudy (anonymous):

agreed